OpenStudy (anonymous):

f(x)=1/4(x+2)^2+4 Vertex is___(4,-2)__ The line of symmetry is_x=-2____ What is the max/min value of f(x)_(4,-2)____ Is this correct?

OpenStudy (anonymous):

no vertex = (-2,4) axis of symmetry: y =-2 min: (-2,4)

OpenStudy (anonymous):

you just switched things is all; no worries

OpenStudy (anonymous):

;-/ I will try another one...hold on just a sec

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

\[y-4=4 \times \frac 4 {16}(x+2)^2 \] Vertex is at \((-2,4)\) Axis of symmetry is \( x=-2 \) Minimum value is at it's vertex \((-2,4)\); There will be no maximum value as this parabola is concave upward.

OpenStudy (anonymous):

No, if it was x = -2, the axis would go through the whole x axis line. the whole point is to divide the parabola

OpenStudy (anonymous):

Thus, : y = (-2,4)

OpenStudy (anonymous):

The axis of symmetry is parallel to the y-axis. \(x=-2 \) is correct.

OpenStudy (anonymous):

no: I think the line proves otherwise...

OpenStudy (anonymous):

f(x)=3x^2-12x+15 Vertex: (2,3) What is the equation of the line of symmetry: 2 What is the max/min:(2,3)

OpenStudy (anonymous):

Either you have a problem with understanding axis of symmetry or you don't know how the line \( x=-2 \) and \(y=-2\) will look.

OpenStudy (anonymous):

@alisa845: Please post it as a new question.-Thanks

OpenStudy (anonymous):

Ok